OpenStudy (anonymous):

the area of the rectangel is 6 square feet. if its width 1 3/7 , fins its length. Any help?

OpenStudy (lalaly):

area of rectangle = length * width so length = area/width

OpenStudy (anonymous):

still doin get it

OpenStudy (lalaly):

length = area/width = 6/(10/7) = 21/5 = 4 1/5

OpenStudy (anonymous):

much appercaited

OpenStudy (lalaly):

anytime :D

OpenStudy (anonymous):

yu got a number lol